OpenStudy (anonymous):

MEDAL AND FAN Jorge plans to paint a bedroom wall that is shaped like a trapezoid. The bottom edge of the wall is 22.5 feet long, and the top edge of the wall is 9.5 feet long. If the wall is 8 feet tall, what is the area of the wall? 1. 256 ft2 2. 38 ft2 3. 90 ft2 4. 128 ft2

OpenStudy (anonymous):

3

OpenStudy (wolf1728):

Trapezoid Area = median * height median = (9.5 + 22.5) / 2 = 16 Area = 16 * 8 = 128 square feet Answer is #4